A triangular region of the community college parking lot was measured. The measure of the first angle of this triangle is triple the measure of the second angle. The measure of the third angle is 6 degrees less than double the measures of the second angle. What are the measurements of all three angles?
Answer by arunpaul(104)

let the angles of a triangle be A,B,C
A = 3B
C=2B-6
again we know A+B+C=180 degree
substituting the value of A and C in the abve equation we have
3B +B +2B-6 =180
or 6B = 180+6= 186
or B = 186/6 =31
so A = 3B = 3*31= 93 degree
C = 2B-6 = 2*31-6 = 62-6=56
so A = 93 degree
B = 31 degree
C = 56 degree
Ans